Use the following data to find the direct labor efficiency variance if the company produced 3,500 units during the period.

Direct labor standard (4 hrs. @ $6.95/hr.) $ 27.80 per unit
Actual hours worked 12,100
Actual rate per hour $ 7.40

Multiple Choice

  • $5,445 unfavorable.

  • $6,300 unfavorable.

  • $6,300 favorable.

  • $13,205 favorable.

  • $5,445 favorable
Answer #1

Labor efficiency variance

= (SH-AH) *SR

= (3500*4 - 12,100)*6.95

= (14,000 - 12,100)*6.95

= 13,205 F